# Random Evolutionary Dynamics in Predator-Prey Systems Yields Large, Clustered Ecosystems

### v1

@@ 4 Formation of genealogical and functional clusters
For the prey species, this function $d$ defines a distance in an $S_{\mathrm{pred}}$-dimensional space, hence a clustering algorithm like kmeans can be used to define clusters in this $S_{\mathrm{pred}}$-dimensional space. For the predator species, the same holds but then with an $S_{\mathrm{prey}}$-dimensional space. We will call this kmeans-distance clustering (KDC) and note that one clustering input parameter is the number of clusters one demands it to find.
@@
The question is now: do these three different clustering approaches lead to similar clusters? We test for similarity as any clustering algorithm is in essence heuristic: different algorithms or similar algorithms with different parameters, result in different clusters. We will first explore this question for a fixed growth rate $r$. Clustering analysis with the aforementioned methods can be done at a given stage of the evolutionary process, but the high and constantly changing dimension of the parameter space makes it difficult to track the emergence of clusters over time. To visualize the emergence of functional clusters, we study the case with evolution in the growth rate $r$ of the prey species in sec. 4.3.
@@ 4.1 Clustering with fixed growth rate: Testing KDC, GDC and IMC algorithms
In Fig. 8, we show the genealogical structures of the predator and prey species for the same parameter setting as the simulations in sec. 3. Only the species alive at the end of the simulation are shown; not the full tree with extinct species. Notice that there are six resp. nine well-defined clusters found by GDC when we take as genealogical distance cut-off $t=10^{6}$. These clusters are persistent in the sense that a large range of cut-off values around $t=10^{6}$ would result in the same clusters. We find the same clusters when we apply KDC to the phylogenetic distance and force it to find six resp. nine clusters, showing the consistency of the two clustering approaches. However, within the KDC we can further examine the similarity within the clusters with the silhouette criterion [21]. Based on this criterion, the optimal number of clusters is five resp. eight: the application of the silhouette criterion merged two clusters into one.
@@ 4.2 Variability in speciation affects clustering
Determining clustering distances with a cut-off is not always as clear as in Fig. 8. Splitting of branches in the range of $[10^{6},1.5\cdot 10^{6}]$ is not as common as outside this region, but certainly happens. The value of $\eta$ is particularly important to consider for the branching. When $\eta$ is decreased and the number of species hence increases (see Fig. 5), the number of branches also increases, which in turn makes it harder to distinguish between different clusters. However, this is not the only factor at play. To examine the ability of the model parameters to change the structure in the genealogical tree, we run simulations twelve times for a range of $\eta$ values and make a normalized histogram of the branching times. As Fig. 9(a) shows, for large $\eta$ values (indicated in red), the majority of the branching times are close to the end of the simulation, indicating that most of the species are in a small number of clusters and many other branches went extinct. On the other hand, for small $\eta$ (indicated in blue), many branches survive as indicated by the high peak at long branching times. To evaluate this observation systematically, we compute the skewness, i.e. a measure of the asymmetry, of the histograms as shown in Fig. 9(a). The result is shown in Fig. 9(b). Indeed, the skewness increases with the variability $\eta$, indicating that the mass of the histograms switches from left to right, i.e. from branching times at the beginning of the simulation to branching times at the end. Hence, the strength of the evolution is shown in the assembly of the community, which is supported by data [31].
@@
The results from Fig. 10 together with the results from Fig. 5 explain our default choice of $\eta=0.02$, as we want rich dynamics, i.e. small $\eta$, but we also want to see cluster formation, i.e. high $\eta$. For values of $\eta$ around $0.02$, the distribution of branching times is fairly symmetric and the minimum of the histogram is around the middle of the time interval $[0,3\cdot 10^{6}]$, indicating that the cluster formation as in Fig. 8 is typical for this $\eta$ value.
@@ 4.3 Robust clustering with evolution in the growth rate
Figure 12: Obtained from the same data as in Fig. 6. Figure (a) shows for each introduced new prey species the time of speciation versus its value of $r_{i}$. Clearly, there is a drift upwards in the values of $r_{i}$. The dots are colored depending on the color of the ancestor. That is, each color is a small perturbation of the RGB color of the ancestor. This highlights the different branches in the evolutionary dynamics. Note however that this figure does not contain any information on the success of each species, i.e. if it survives and for how long. Figure (b) shows the value of $r_{i}$ versus the predation, i.e. the corresponding row sum of $P$, for the 38 alive prey species at the end of the simulation. The same coloring as in Fig. (a) is used. The two ellipses indicate the lower branch (left) and upper branch (right) of Fig. (a).
@@ 4.3.1 Correspondence of functional and interaction clustering
Figs. 11(a) and 11(b) show that when the green-colored prey species from the lower branch overlap with the red-colored prey species from the upper branch in the $r_{i}$ versus time plot, their rows of $P$ are not necessarily similar as indicated by the two different ellipses. Therefore, the species in the upper and lower branches are not just different in the growth rate $r_{i}$, but also differ in their functional interactions with the predators. To further investigate this, we apply the same IMC algorithm as used before in the fixed growth rate context to the interaction matrix $A$. As Fig. 13 shows, the $2\times 2$ (predators versus prey) block structure can be divided into a smaller $4\times 4$ structure (two groups of predators versus two groups of prey) such that the interaction within these blocks is significantly stronger than the interaction between the different blocks. This shows that there are indeed two main clusters of interacting prey and predator species. Importantly, the two clusters found by the IMC algorithm exactly overlap with the upper and lower branch of the evolutionary tree. Hence, species with a strong predator-prey interaction share an evolutionary history, which is strong evidence of co-evolution in our model.
@@ 5 Discussion & Outlook
The fact that structure in ecosystems is not random but clustered is an important observable to reproduce while modeling evolutionary dynamics [31]. So far, the approaches for resource-consumer models [5, 10], where consumption is normalized, lead to a small number of clusters determined by the normalization. Alternatively, the competitive LV approach in [35] leads to more clusters, but these are largely formed by mother-daughter interactions and therefore not reflective of real-world ecosystems [2, 25]. Enforcing two trophic layers with predator-prey interactions resulted in stable evolutionary dynamics as shown in sec. 3. In the case of large and intermediate variability, clusters of predators and prey arose which are robust against variation in the growth rate, but in the limit of small variation, there are too many branches in the evolutionary tree to meaningfully define clusters.
@@
The speciating LV model with two trophic layers as discussed here has limitations. In real-world ecosystems, all three types of interaction, competition, predation and mutualism, are possible instead of just predation. We also investigated ecosystem modeling without enforcing predator-prey relations, i.e. allowing competition and mutually beneficial relations without specified trophic layers. In particular, in Appendix A, we study a generalized LV equation with speciation without restrictions on the interaction matrix, but this approach did not lead to stable evolutionary dynamics.
In the current approach, resources are not explicitly modeled like, for example, in [5]. Hence, the prey species do not compete for resources and each prey species follows its own logistic growth rate, independent of other prey species. Combining predator-prey interactions with resource-consumer models, hence creating three trophic layers, will be an important next step in extending the evolutionary dynamics.
In our speciating predator-prey LV model, the interaction matrix $A$ is enforced to be anti-symmetric (up to a factor $\beta$). This $\beta$ is fixed for all predator and prey combinations and therefore leaves no room for specialization, as a predator can consume each prey species with similar effort and reward. Letting the $\beta$ be species-dependent and vary with the speciation process allows specialization among the predators. However, in the current setup, this would imply that predators with higher $\sum_{j}\beta_{ij}$ have a significant advantage over their competitors leading to so-called “Darwinian Demons”, so a budget for the vector $\beta$ would be necessary in this case.
All the species in our model are to some extent artificially similar, as the values of $\delta,K$, and $\beta$ are fixed. The carrying capacity $K$ can be removed by explicitly introducing resources and $\beta$ has been discussed above, but for the mortality $\delta$, it is not a priori clear how to introduce stochastic variability. We have to assume that $\delta>0$, but introducing noise into this parameter automatically leads to a downward drift as a lower mortality rate is advantageous. If a similar approach is taken as in sec. 4, there will be species with negative $\delta$ on the timescale of the simulation (result not shown). Hence, a more detailed description of mortality is needed before noise can be added to it; this is left as future work.
@@ 6 Conclusions
We studied random speciation in a predator-prey LV model and found that this stochastic approach of adding new species leads to large complex ecosystems that (a) are stable, (b) have their ecological and evolutionary timescales intertwined and (c) show clustering of species, both functional and genealogical. The dynamics observed is robust to various model parameter variations and cluster definitions. Our results indicate that adding stochastic speciation to a coupled ODE system has both mathematically novel and biologically relevant implications for the better modeling of ecosystems.
@@ 7 Acknowledgments
The work of CH is funded by the Dutch Institute for Emergent Phenomena (DIEP) at the University of Amsterdam via the program Foundations and Applications of Emergence (FAEME).
@@ Appendix A Generalized Lotka-Volterra equations with speciation
In the main text, we enforced the model to have two trophic layers, i.e. Eq. (2) and (3). In this appendix, we showcase the necessity of this two trophic layer condition for the current setup with speciation. To this end, we relax this condition and, following the approach of [35], we start with Generalized Lotka-Volterra (GLV) equations
@@
The biomass $x_{i}$ of each of the $N$ species present in the system follows a logistic model with growth rate $r_{i}>0$ and carrying capacity $K$, on top of which interaction with different species is possible. In the community matrix $A$, each component $A_{ij}$ describes the influence species $j$ has on $i$. In general, $A_{ij}\neq A_{ji}$, so $A$ is not symmetric. Furthermore, we take $A_{ii}=0$, as the self-interaction is modeled by the logistic term. For simplicity, we fix the carrying capacity $K$. The signs of $A_{ij}$ and $A_{ji}$ describe the relationship between the species $i$ and $j$. Note that the signs of $A_{ij}$ are not fixed as, in contrast to the main text, there are no pre-described trophic layers.
@@
where $\psi_{c}$ is a perturbation of the column $A_{*,i}$ and $\psi_{r}$ is a perturbation of the row $A_{i,*}$. In this formulation of speciation, the interaction between daughter and ancestor is given by $\tilde{A}_{(n+1)i}=A_{ii}+\psi_{r}(i)=\psi_{r}(i)\sim\mathcal{N}(0,\eta^{2})$, which is weak as the standard deviation $\eta$ is small. That is, mother and daughter are very similar. However, we expect there to be strong competition between the daughter and ancestor species as their niches are similar. Therefore, to model the competition between these two species, we take $\tilde{A}_{(n+1)i}=-\gamma+\psi_{r}(i)$ and for $\tilde{A}_{i(n+1)}$ we take the same approach. For the upcoming figures, we use $\gamma=0.4$.
@@
In contrast to the two trophic layer model, the speciating GLV model has no built-in feedback mechanism to bound the biomass. Each species has a carrying capacity of $K$, but when $A$ has enough positive values, this self-limiting term can be overcome and the biomass in the system diverges. To investigate if, and when, this happens, we simulate the system and track the biomass $B$, the species abundance $S$ (as defined in the main text), and the average $\bar{A}$ of the interaction matrix A. We expect that the biomass initially will remain below $KS$. However, as Fig. A.1 shows, the moment $\bar{A}$ becomes positive, the biomass grows unboundedly and thus unrealistically.
@@
Of course, the observation “a positive value of $\bar{A}$ leads to divergence” is not a mathematical theorem. Indeed, there are $2\times 2$ matrices with positive averages that lead to stable dynamics. Furthermore, the exact timing of the divergence also depends on the value of $\gamma$. When the mother-daughter interaction becomes stronger, there can be a delay in the divergence of the biomass (results not shown). However, we have not been able to find a parameter setting in which the unbounded growth of the biomass in the speciating GLV model did not occur.
@@ A.1 Battling biomass divergence
The observation of biomass divergence is remarkably robust. Fine-tuning the several parameters in the speciating GLV model appears to only slow down the moment of biomass divergence, but cannot prevent this unphysical behavior. This unbounded growth is possible because the resources are not explicitly modeled. In a resource-consumer model like the MacArthur model, considered in a similar context as in this article in [5], divergence is not possible. We will now briefly discuss a possible solution to stabilize the unphysical ecosystem dynamics.
One way to prevent divergence is to enforce that the coefficients in the interaction matrix are nonpositive. This can be realized by setting all positive values in $A$ to zero after a speciation step. However, the drift upwards in the coefficients of $A$ due to speciation is still observed (result not shown). This will result in a situation with many species, but the competition is effectively only with recent ancestors, as in every speciation step a $-\gamma$ niche interaction term is introduced. To show this effect, we follow [35] and cluster the living species at the end of a simulation according to their interaction strengths, using the same ICM algorithm [7] as in the main text. Figs. 1(a) and 1(b) show the unclustered and clustered interaction matrix $A$. These figures indeed show that the surviving 203 species are clustered in five groups of species that strongly interact. The question is now if this clustering correlates to genealogical distance. The matrix of all genealogical distances is shown in Fig. 1(c), but when we sort it according to the clustering in Fig. 1(b), we see in Fig. 1(d) that the genealogical distance within these clusters is short. In terms of the model, this implies that the interaction is dominated by the introduction of the mother-daughter interaction $\gamma$. As pointed out in [35], this outcome supports Darwin’s competition-relatedness hypothesis. However, studies in real-world ecosystems suggest that competition is much more based on functional similarities instead of genealogical similarities [2, 25]. Hence, the absence of interaction between the different clusters here appears to be artificial and not reflective of real ecosystems. In contrast, this interaction between clusters is present in the two trophic layer model as discussed in the main text, see in particular Fig. 13.

### v2

@@ 4 Formation of genealogical and functional clusters
For the prey species, this function $d$ defines a distance in an $S_{\mathrm{pred}}$-dimensional space, hence a clustering algorithm like kmeans can be used to define clusters in this $S_{\mathrm{pred}}$-dimensional space. For the predator species, the same holds but then with an $S_{\mathrm{prey}}$-dimensional space. We will call this kmeans-distance clustering (KDC) and note that one clustering input parameter is the number of clusters one demands it to find. See Appendix B for a discussion on the different settings of this algorithm.
@@
Does the choice of clustering algorithm affect the clustering observation? We test for similarity as any clustering algorithm is in essence heuristic: different algorithms or similar algorithms with different parameters, result in different clusters. We will first explore this question for a fixed growth rate $r$. Clustering analysis with the aforementioned methods can be done at a given stage of the evolutionary process, but the high and constantly changing dimension of the parameter space makes it difficult to track the emergence of clusters over time. To visualize the emergence of functional clusters, we study the case with evolution in the growth rate $r$ of the prey species in sec. 4.3.
@@ 4.1 Clustering with fixed growth rate: Testing KDC, GDC and IMC algorithms
In Fig. 9, we show the genealogical structures of the predator and prey species for the same parameter setting as the simulations in sec. 3. Only the species alive at the end of the simulation are shown; not the full tree with extinct species. Notice that there are six respectively nine well-defined clusters found by GDC when we take as genealogical distance cut-off $t=10^{6}$. These clusters are persistent; a large range of different cut-off values around $t=10^{6}$ would result in the same clusters. We find the same clusters when we apply KDC to the phylogenetic distance and force it to find six respectively nine clusters, showing the consistency of the two clustering approaches. However, with the KDC, we can further examine the similarity within the clusters with the silhouette criterion [22]. Based on this criterion, the optimal number of clusters is five respectively eight: the application of the silhouette criterion merged two clusters into one.
@@ 4.2 Variability in speciation affects clustering
Determining clustering distances with a cut-off is not always as clear as in Fig. 9. Splitting of branches in the range of $[10^{6},1.5\cdot 10^{6}]$ is not as common as outside this region, but certainly happens. The value of $\eta$ is particularly important to consider for the branching. When $\eta$ is decreased and the number of species hence increases (see Fig. 6), the number of branches also increases, which in turn makes it harder to distinguish between different clusters. However, this is not the only factor at play. To examine the ability of the model parameters to change the structure in the genealogical tree, we run simulations twelve times for a range of $\eta$ values and make a normalized histogram of the branching times. On the one hand, as Fig. 11(a) shows, for large $\eta$ values (indicated in red), the majority of the branching times are close to the end of the simulation, indicating that most of the species are in a small number of clusters and many other branches went extinct. On the other hand, for small $\eta$ (indicated in blue), many branches survive as indicated by the high peak at long branching times. To evaluate this observation systematically, we compute the skewness, i.e. a measure of the asymmetry, of the histograms as shown in Fig. 11(a). The result is shown in Fig. 11(b). Indeed, the skewness increases with the variability $\eta$, indicating that the mass of the histograms switches from left to right, i.e. from branching times at the beginning of the simulation to branching times at the end. Hence, the strength of the evolution is shown in the assembly of the community, which is supported by data [32].
@@
The results from Fig. 11 together with the results from Fig. 6 explain our default choice of $\eta=0.02$. To study branching and clustering in a stochastic model we need sufficient statistics for both. We thus aim to be in an intermediate noise level as we need many speciation events, i.e. small $\eta$. However, we also need sufficient statistics on the clustering observations, i.e. high $\eta$. We found that for values of $\eta$ around $0.02$, the distribution of branching times is fairly symmetric and the minimum of the histogram is around the middle of the time interval $[0,3\cdot 10^{6}]$, indicating that the cluster formation as in Fig. 9 is typical for this $\eta$ value.
@@ 4.3 Robust clustering with evolution in the growth rate
Figure 13: Obtained from the same data as in Fig. 7. Figure (a) shows for each introduced new prey species the time of speciation versus its value of $r_{i}$. Clearly, there is a drift upwards in the values of $r_{i}$. The dots are colored depending on the color of the ancestor. That is, each color is a small perturbation of the RGB color of the ancestor. This highlights the different branches in the evolutionary dynamics. Note however that this figure does not contain any information on the success of each species, i.e. whether it survives and for how long. Figure (b) shows the value of $r_{i}$ versus the predation, i.e. the corresponding row sum of $P$, for the 38 alive prey species at the end of the simulation. The same coloring as in Fig. (a) is used. The two ellipses indicate the lower branch (left) and upper branch (right) of Fig. (a).
@@ 4.3.1 Correspondence of functional and interaction clustering
Figs. 13(a) and 13(b) show that when the green-colored prey species from the lower branch overlap with the red-colored prey species from the upper branch in the $r_{i}$ versus time plot, their rows of $P$ are not necessarily similar as indicated by the two different ellipses. Therefore, the species in the upper and lower branches are not just different in the growth rate $r_{i}$, but also differ in their functional interactions with the predators. To further investigate this, we apply the same IMC algorithm as used before in the fixed growth rate context to the interaction matrix $A$. As Fig. 14 shows, the $2\times 2$ (predators versus prey) block structure can be divided into a smaller $4\times 4$ structure (two groups of predators versus two groups of prey) such that the interaction within these blocks is significantly stronger than the interaction between the different blocks. This shows that there are indeed two main clusters of interacting prey and predator species. Importantly, the two clusters found by the IMC algorithm exactly overlap with the upper and lower branch of the evolutionary tree. Hence, species with a strong predator-prey interaction share an evolutionary history, which is strong evidence of co-evolution in our model.
@@ 5 Discussion & Outlook
The fact that structure in ecosystems is not random but clustered is an important observable to reproduce while modeling evolutionary dynamics [32]. So far, the approaches for resource-consumer models [5, 9], where consumption is normalized, lead to a small number of clusters determined by the normalization. Alternatively, the competitive LV approach in [39] leads to more clusters, but these are largely formed by mother-daughter interactions and therefore not reflective of real-world ecosystems [2, 26]. Enforcing two trophic layers with predator-prey interactions resulted in stable evolutionary dynamics as shown in sec. 3. In the case of large and intermediate variability, clusters of predators and prey arose, and the clusters are robust against variation in the growth rate, but in the limit of small variation, there are too many branches in the evolutionary tree to meaningfully define clusters.
@@
The speciating LV model with two trophic layers as discussed here has limitations. In real-world ecosystems, all three types of interaction, competition, predation and mutualism, are possible instead of just predation. We also investigated ecosystem modeling without enforcing predator-prey relations, i.e. allowing competition and mutually beneficial relations without specified trophic layers. In particular, in Appendix A, we study a generalized LV equation with speciation without restrictions on the interaction matrix, but this approach consistently showed a divergent biomass and hence did not lead to stable evolutionary dynamics, rendering cluster analysis moot.
In the current approach, resources are not explicitly modeled like, for example, in [5]. Hence, the prey species do not compete for resources and each prey species follows its own logistic growth rate, independent of other prey species, which in theory allows for unlimited growth of biomass. Combining predator-prey interactions with resource-consumer models, hence creating three trophic layers, will be an important next step in extending the evolutionary dynamics.
In our speciating predator-prey LV model, the interaction matrix $A$ is enforced to be anti-symmetric, up to a factor $\beta$. This $\beta$ is fixed for all predator and prey combinations and therefore leaves no room for specialization, as a predator can consume each prey species with similar effort and reward. Letting the $\beta$ be species-dependent and vary with the speciation process allows specialization among the predators. However, in the current setup, this would imply that predators with higher $\sum_{j}\beta_{ij}$ have a significant advantage over their competitors leading to so-called “Darwinian Demons”, so a budget for the vector $\beta$ would be necessary in this case, for example by fixing its norm.
All the species in our model are to some extent artificially similar, as the values of $\delta,K$, and $\beta$ are fixed. The carrying capacity $K$ can be removed by explicitly introducing resources and $\beta$ has been discussed above, but for the mortality $\delta$, it is not a priori clear how to introduce stochastic variability. We have to assume that $\delta>0$, but introducing noise into this parameter automatically leads to a downward drift as a lower mortality rate is advantageous. If a similar approach is taken as in sec. 4, there will be species with negative $\delta$ on the timescale of the simulation (result not shown). Hence, a more detailed description of mortality is needed before it can be modified by noise; this is left as future work.
@@ 6 Conclusions
We studied random speciation in a predator-prey LV model and found that this stochastic approach of adding new species leads to large complex ecosystems that (a) are stable, (b) have their ecological and evolutionary timescales intertwined and (c) show clustering of species, both functional and genealogical. The observed dynamics are robust to various model parameter variations and cluster definitions. Our results indicate that adding stochastic speciation to a coupled ODE system has both mathematically novel and biologically relevant implications for the better modeling of ecosystems.
@@ 7 Acknowledgments
The work of CH is funded by the Dutch Institute for Emergent Phenomena (DIEP) at the University of Amsterdam via the program Foundations and Applications of Emergence (FAEME). We thank Marco Saltini for constructive remarks on earlier versions of this paper.
@@ Appendix A Generalized Lotka-Volterra equations with speciation
In the main text, we enforced the model to have two trophic layers, i.e. Eq. (2) and (3). In this appendix, we argue for the necessity of this two-trophic layer condition for the current setup with speciation, by showing that the more general case leads to unphysical behavior. We relax the tropic layer condition, following the approach of [39], we start with Generalized Lotka-Volterra (GLV) equations
@@
The biomass $x_{i}$ of each of the $N$ species present in the system follows a logistic model with growth rate $r_{i}>0$ and carrying capacity $K$, on top of which interaction with different species is possible. In the community matrix $A$, each component $A_{ij}$ describes the influence species $j$ has on $i$. In general, $A_{ij}\neq A_{ji}$, so $A$ is not symmetric. Furthermore, we take $A_{ii}=0$, as the self-interaction is modeled by the logistic term. For simplicity, we fix the carrying capacity $K$. The signs of $A_{ij}$ and $A_{ji}$ describe the relationship between the species $i$ and $j$. Note that the signs of $A_{ij}$ are not fixed as, in contrast to the main text, there are no pre-described trophic layers. In contrast to the two trophic layer model, the speciating GLV model therefore has no built-in feedback mechanism to bound the biomass. Each species has a carrying capacity of $K$, but when $A$ has enough positive values, this self-limiting term can in principle be overcome and the biomass in the system diverges. The question is whether the stochastically generated $A_{ij}$ may allow for a stable solution at long timescales. As we will see below, we have numerically not been able to find a set of parameters that keeps the biomass finite, but we have no proof that such a divergence must always occur in this model implementation.
@@
where $\psi_{c}$ is a perturbation of the column $A_{*,i}$ and $\psi_{r}$ is a perturbation of the row $A_{i,*}$. In this formulation of speciation, the interaction between daughter and ancestor is given by $\tilde{A}_{(n+1)i}=A_{ii}+\psi_{r}(i)=\psi_{r}(i)\sim\mathcal{N}(0,\eta^{2})$, which is weak as the standard deviation $\eta$ is small. That is, mother and daughter are very similar. However, we expect there to be strong competition between the daughter and ancestor species as their niches are similar. Therefore, to model the competition between these two species, we take $\tilde{A}_{(n+1)i}=-\gamma+\psi_{r}(i)$ and for $\tilde{A}_{i(n+1)}$ we take the same approach. For the figures below, we use $\gamma=0.4$.
@@
Biomass divergence in Eq. A.1 occurs when the $A_{ij}$ become positive enough on average to induce an unbounded growth for some $x_{i}$. To investigate if, and when, this happens during speciation, we simulate the system and track the biomass $B$, the species abundance $S$ (as defined in the main text), and the average $\bar{A}$ of all the interaction matrix coefficients $A_{ij}$. We expect that the biomass initially will remain below $KS$ but that an increase in $\bar{A}$ should eventually precipitate in a biomass divergence. Indeed, as Fig. A.1 shows, the moment $\bar{A}$ becomes positive, the biomass grows without bound.
@@
We cannot prove under which conditions for $\bar{A}$ the biomass divergence occurs. Indeed, there are counterexamples of $2\times 2$ matrices with positive averages that lead to stable dynamics. The general case of $N\times N$ matrices is however not easily considered. Stability of biomass also has a temporal dimension: the exact timing of the divergence depends on the value of $\gamma$: we have observed that when the mother-daughter interaction becomes stronger, the divergence of the biomass typically occurs at later times (results not shown). While absence of proof is not proof of absence, we have not been able to find any parameter setting in which the biomass in the speciating GLV model remains finite. The observation of biomass divergence is in fact remarkably robust. Empirical evidence suggests that adjusting the several parameters in the speciating GLV model can only slow down the moment of biomass divergence, but cannot prevent this unphysical behavior.
@@ A.1 Battling biomass divergence
We will now briefly discuss a possible more biologically motivated or ad-hoc solution to stabilize the unphysical ecosystem dynamics. One way to constrain species growth is by coupling the growth rate to a finite (replenishable) resource bath. In a resource-consumer model like the MacArthur model, considered in a similar context as in this article in [5], divergence is not possible because biomass increase depletes resources and suppresses growth rates.
Another way to prevent biomass divergence is to enforce that the coefficients in the interaction matrix are nonpositive. This can be realized by setting all positive values in $A$ to zero after a speciation step. However, the drift upwards in the coefficients of $A$ due to speciation is still observed (result not shown). This will result in a situation with many species, but the competition is effectively only with recent ancestors, as in every speciation step a $-\gamma$ niche interaction term is introduced. To show this effect, we follow [39] and cluster the living species at the end of a simulation according to their interaction strengths, using the same ICM algorithm [7] as in the main text. Figs. 2(a) and 2(b) show the unclustered and clustered interaction matrix $A$. These figures indeed show that the surviving 203 species are clustered in five groups of species that strongly interact. The question now is whether this clustering correlates to genealogical distance. The matrix of all genealogical distances is shown in Fig. 2(c), but when we sort it according to the clustering in Fig. 2(b), we see in Fig. 2(d) that the genealogical distance within these clusters is short. In terms of the model, this implies that the interaction is dominated by the introduction of the mother-daughter interaction $\gamma$. As pointed out in [39], this outcome supports Darwin’s competition-relatedness hypothesis. However, studies in real-world ecosystems suggest that competition is much more based on functional similarities instead of genealogical similarities [2, 26]. Hence, the absence of interaction between the different clusters here appears to be artificial and not reflective of real ecosystems. In contrast, this interaction between clusters is present in the two trophic layer model as discussed in the main text, see in particular Fig. 14.
@@
Appendix B Sensitivity analysis of model parameters
In this appendix, we discuss some of the system parameters from the main text and explain why certain choices were made, and why some parameter choices do not significantly affect the modeling outcomes.
Carrying capacity $K$
We chose to work with a carrying capacity $K=10$. Before we discuss this specific choice, let us study the case of $K\rightarrow\infty$, i.e. no carrying capacity. This choice makes the model linear, and all else being equal, this results in oscillatory dynamics instead of an asymptotic fixed point. This in turn implies that the total biomass fluctuates strongly. An example run is shown in Fig. 1(a). We observe that in this setting, the number of species is small. Furthermore, when speciation is stopped at the end of the simulation, the resulting system with two predator and two prey species is chaotic, see Fig. 1(b). We conclude that a carrying capacity is needed to obtain systems with a high number of species.
Figure B.1: Single run of the model for $K\rightarrow\infty$, i.e. no carrying capacity. As is shown, the total number of predators (red) and prey (blue) $S$ remains low. Out of the six species that remain at $t=5\cdot 10^{5}$, four survive after speciation is stopped. In Figure (b), we show the dynamics of these four remaining species after a long integration time to remove any transients, but the dynamics remains chaotic, as it scored a 0.9987 using the 0-1 chaos test from [15].
For the fixed point (5) to exist, we need $K>\delta/(\beta P)$. The latter equals four for the values in Table 1. In Fig. 2(a), we plot the average number of species $\langle S\rangle$ for a range of $K$-values. Clearly, $\langle S\rangle\sim K^{-1}$, as indicated by the red line which is a fit to $\langle S\rangle=aK^{-1}$ with $a\approx 829$. A fit with a free exponent yields similar results; we choose not to fit an offset because of the limiting behavior of $K\to\infty$ discussed above. The interpretation of the $1/K$ scaling relationship is beyond the scope of this work, but we can see that our model choice of $K=10$ can be considered a trade-off between making sure the model is not very sensitive to $K$, while keeping enough species for the clustering analysis.
Figure B.2: Figure (a) shows $\langle S\rangle$ as a function of the carrying capacity $K$, the red line is a fit to the power law $\langle S\rangle=aK^{-1}$, with $a\approx 829$. Figure (b) shows $\langle S\rangle$ as a function of $f$, the fraction of biomass that is transferred to the daughter species. This figure shows that $f$ does not meaningfully influence the average number of species.
Ratio Ancestor-Daughter species $f$
In the main text, we set the fraction $f$ of biomass that is transferred from the ancestor to the daughter species to 0.05. Changing this parameter does not essentially influence the dynamics, as is shown in Fig. 2(b). Only when $f$ is chosen close to or below the extinction threshold, $\langle S\rangle$ trivially decreases.
Existence Threshold —
The average number of species $\langle S\rangle$ as a function of the extinction threshold has two trivial limits. When the threshold is larger than the total biomass in the system, the number of species is 0. When there is no extinction threshold, no species go extinct as species decay exponentially to 0. Given these extreme limits, we plot the dependence on the threshold over two orders of magnitude and indeed observe the expected negative relation, but the decay is slow, the fit in Fig. 3(b) indicates a weak power law decay with exponent -0.11. Hence, the sensitivity of $\langle S\rangle$ on the extinction threshold is small.
Figure B.3: Figure (a) shows the average number of species $\langle S\rangle$ as a function of the extinction threshold. From this figure, the decay rate is difficult to probe, but in the log-log plot in Figure (b), there is a clear indication of a power law. However, the decay is small as the red line has a slope of $-0.11$.
B.1 Parameters in clustering algorithms
For the genealogical-distance clustering, the dependence of the number of clusters on the cut-off in the genealogical distance can be understood visually from Fig. 9. A smaller distance implies more clusters. For Matlab’s kmeans algorithm, however, it is not immediately clear how the different options influence the outcome. Given the matrix $P$ as defined in (4) and the vector $r$ of growth rates, we write in Matlab
Clusters=kmeans([3*r,P],6,’Distance’,’sqEuclidean’,’Replicates’,500);
silh=silhouette([3*r,P],Clusters,’sqEuclidean’);
mean(silh)
The score for the clustering is given by mean(silh), where a higher score is better. The number of replicates here must be high to ensure repeatability. For a low number of replicates, the clustering is different each turn. Matlab has a range of choices for the distance function, sqEuclidean, cityblock, cosine and correlation. When these different options are applied to the data from Fig. 15 for cluster numbers 5, 6 and 7, all options return the same clustering with 6 clusters as optimal.